OpenStudy (grandmoffchris):

28. A carpenter cut a board into 3 pieces of equal length. One of the pieces was 15.2 cm long. What was the length of the board?

OpenStudy (grandmoffchris):

@pooja195

OpenStudy (aihberkhan):

Let x = smaller piece 5x= larger piece x+5x=72 6x=72 x=12 cm smaller piece 5x=5(12)=60 cm larger piece Check: 12 +60=72 It works!

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @GrandMoffChris

OpenStudy (grandmoffchris):

Thanks

OpenStudy (grandmoffchris):

So umm

OpenStudy (grandmoffchris):

Mrs. Williams has 24 students in her class this year. The number of students she has this year is 120% of the number she had last year. How many students were in Mrs. William's Class last year?

OpenStudy (grandmoffchris):

@pooja195 Halp

OpenStudy (grandmoffchris):

@pooja195 !!!!

OpenStudy (grandmoffchris):

@uybuyvf @RCCB

OpenStudy (grandmoffchris):

HELP ME

OpenStudy (grandmoffchris):

JUST HELP MEEEEEEE

OpenStudy (grandmoffchris):

will someone help or not?

OpenStudy (mrnood):

what does 120% mean as a fraction? did she have MORE or fewer students?

OpenStudy (grandmoffchris):

just read,

OpenStudy (mrnood):

I have read it (and the answer is trivial) - I am Asking YOU to answer the question what is 120% AS A FRACTION

OpenStudy (grandmoffchris):

I HAVE NO IDEA, I DONT KNOW FRACTIONSA

OpenStudy (mrnood):

so how do you expect to answer this question then?

OpenStudy (grandmoffchris):

thats why im ASKING IT

OpenStudy (grandmoffchris):

THATS WHAT OPEN STUDY IS FOR

OpenStudy (grandmoffchris):

THEY HELP ME RESOLVE IT

OpenStudy (mrnood):

it is NOT for being given the answer you need to have the basic knowledge to tackle a question if I give oyu the answer you learn NOTHING

OpenStudy (grandmoffchris):

TAHTS WHY I SAID EXPLAIN TO ME

OpenStudy (grandmoffchris):

OR RESOLVE

OpenStudy (mrnood):

120% means 120/100 so THAT is th fraction more that she has (please don't shout)

OpenStudy (grandmoffchris):

THEY HELP ME RESOLVE IT

OpenStudy (anonymous):

xD please do shout Grand show who ever this is ...

OpenStudy (mrnood):

so the number she has this year = 120/100 times the number she had last year

OpenStudy (grandmoffchris):

so tf that means?

OpenStudy (mrnood):

that means I have given you the whole information you need to answer the question however - your attitude tells me that you have no intention of learning HOW to do this

OpenStudy (grandmoffchris):

know wht ? shut up and gtfo

OpenStudy (grandmoffchris):

pooja teaches me better

OpenStudy (mrnood):

the number she has this year = 120/100 times the number she had last year 24 = 120/100 x (the answer)

OpenStudy (anonymous):

xDDDDDDDDD -dies-